### v1

# Plan2Align: Predictive Planning Based Test-Time Preference Alignment in Paragraph-Level Machine Translation

Abstract
Machine Translation (MT) has been predominantly designed for sentence-level translation using transformer-based architectures. While next-token prediction based Large Language Models (LLMs) demonstrate strong capabilities in long-text translation, non-extensive language models often suffer from omissions and semantic inconsistencies when processing paragraphs. Existing preference alignment methods improve sentence-level translation but fail to ensure coherence over extended contexts due to the myopic nature of next-token generation. We introduce Plan2Align, a test-time alignment framework that treats translation as a predictive planning problem, adapting Model Predictive Control to iteratively refine translation outputs. Experiments on WMT24 Discourse-Level Literary Translation show that Plan2Align significantly improves paragraph-level translation, achieving performance surpassing or on par with the existing training-time and test-time alignment methods on LLaMA-3.1 8B.
Plan2Align: Predictive Planning Based Test-Time Preference Alignment
in Paragraph-Level Machine Translation
Kuang-Da Wang1,† Teng-Ruei Chen1,† Yu Heng Hung1 Shuoyang Ding2 Yueh-Hua Wu2 Yu-Chiang Frank Wang2Chao-Han Huck Yang2Wen-Chih Peng1Ping-Chun Hsieh1 1National Yang Ming Chiao Tung University  2NVIDIA {gdwang.cs10@,raychen.cs12@,wcpeng@cs.,pinghsieh@}nycu.edu.tw{hucky,krisw,shuoyangd}@nvidia.com equal contribution
1 Introduction
Machine Translation (MT) has primarily leveraged transformer-based encoder-decoder neural architectures, as exemplified by models like M2M-100 (Fan et al., 2021) and MT5 (Xue et al., 2021). However, the emergence of Large Language Models (LLMs), such as the GPT series (Brown et al., 2020; Achiam et al., 2023), LLaMAs (Touvron et al., 2023a, b), and Gemma (Team et al., 2024), has demonstrated remarkable efficacy in a wide range of NLP tasks (Stiennon et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2023; Yu et al., 2024; Zhong et al., 2024), including multilingual translation. This progress has led to increasing interest in leveraging LLMs for MT, particularly for handling long-form text. LLMs excel at long-text translation due to their ability to model dependencies across extended contexts, enabling improved coherence and fluency. Models such as Tower (Alves et al., 2024) and GPT-4o (OpenAI, 2024) have demonstrated state-of-the-art performance, surpassing traditional MT systems in both fluency and contextual consistency.
Figure 1: Machine translation via predictive planning from the perspective of trajectory optimization.
However, non-extensive (i.e., smaller than 10B) LLMs often struggle with long-text translation, exhibiting sentence omissions, semantic inconsistencies, and hallucinations (Wu et al., 2024). These limitations stem from the myopic nature of next-token prediction, where local fluency is prioritized at the cost of global coherence, leading to degraded translation quality in extended discourse. One-pass translation remains inherently limited, as it lacks iterative refinement, making it prone to over-translation, under-translation, and compounding errors (Mandya et al., 2020). While traditional MT systems incorporate techniques such as re-ranking, iterative decoding, and constrained beam search to improve coherence and fidelity, LLM-based MT operates purely in a next-token prediction paradigm, lacking an explicit self-correction mechanism (Hadj Ameur et al., 2019; Feng et al., 2024). As a result, errors introduced early in the generation process propagate across extended text spans, leading to semantic drift, loss of contextual dependencies, and degraded translation quality in long-form content.
Recent research has investigated preference alignment as a means to improve translation quality in small-scale LLMs, leveraging preference data to guide optimization. Methods such as CPO (Xu et al., 2024b) and WAP (Wu et al., 2024) have shown promising results in aligning translations to preferences. However, existing methods face two fundamental challenges in long-text translation:
  1. 1.
    Preference alignment techniques operate at specific granularity levels; when applied at the long-text level, they often struggle to preserve sentence coherence across paragraphs.
  2. 2.
    Unlike training-time alignment, which requires costly fine-tuning, testing-time alignment methods must incorporate a technique to recover from past errors or low-quality answers caused by compounding errors during inference. This is crucial because these past mistakes can lead to further errors, as earlier responses influence the generation of answers to subsequent questions (Mandya et al., 2020).
To address these challenges, we propose Plan2Align, a test-time alignment method that employs a self-rewriting framework based on the principle of Model Predictive Control (MPC), a classic predictive planning method for trajectory optimization (Rao, 2014) widely used in robotics (Tassa et al., 2012, 2014; Nagabandi et al., 2020) and motion planning (Ji et al., 2016). Rather than fine-tuning LLMs, Plan2Align iteratively refines translations by integrating high-quality segments from previous iterations at test time, enhancing coherence and fluency in long-text translation.
This study focuses on paragraph-level MT, where translation quality is evaluated beyond isolated sentences. Specifically, we define context-level as a group of adjacent, semantically related sentences, while a paragraph consists of multiple interconnected contexts. Ensuring high-quality translation at the context level is crucial for maintaining coherence and accuracy at the paragraph level. Specifically, this study is twofold: (i) to develop a test-time alignment method that matches or surpasses fine-tuning approaches, and (ii) to enhance paragraph-level translation by improving consistency and fidelity across context-level translations.
As a predictive planning method, MPC optimizes a sequence of decisions by continuously adjusting future actions based on past observations (Camacho and Bordons, 2007; Kouvaritakis and Cannon, 2016). Inspired by this, we draw an analogy between trajectory optimization and translation optimization, where prior translations inform subsequent refinements, and recast MT as an iterative optimization process rather than a one-pass mapping. Figure 1 illustrates this correspondence, highlighting the intuition behind our approach. While LLMs excel in contextual understanding, their one-pass translations remain suboptimal. By iteratively leveraging high-quality past translations, our method enhances coherence and discourse integrity, improving overall quality.
However, directly applying MPC to paragraph-level MT is impractical when entire paragraph-level translation histories are fed into small-scale LLMs due to their limited contextual capacity. To adapt MPC to paragraph-level MT, we introduce two novel mechanisms: (i) Model-predictive context alignment: We modify MPC to selectively retain high-quality contexts from multiple paragraph-level translations with the help of a context buffer, instead of relying on a single best trajectory. This ensures effective accumulation and utilization of translation experience for aligning source and translation contexts. (ii) Self-rewriting tasks: We redefine MT as a self-rewriting task, a structured prompting mechanism that enhances discourse coherence, enables each iteration to build upon prior high-quality translations, and ultimately improves fluency and consistency. By integrating these two strategies, Plan2Align achieves efficient translation refinement within a limited number of iterations.
To validate our approach, we apply Plan2Align to LLaMA-3.1 8B on the WMT’24 Discourse-Level Literary Translation Benchmark, using 6K paragraph-level preference data for test-time alignment. Our method demonstrates substantial improvements, achieving performance on par with or surpassing training-time alignment methods while outperforming existing test-time alignment approaches. Plan2Align bridges the gap in LLM-based paragraph-level translation, introducing a novel test-time alignment framework that ensures high translation quality while maintaining context-level consistency. Moreover, our test-time alignment method can augment other fine-tuning approaches, leveraging a small number of iterations at test time to achieve even better translation performance.
2 Related Work
2.1 Document-Level Machine Translation
Document-level neural machine translation can be broadly divided into two categories: the sentence-to-sentence (sen2sen) approach and the document-to-document (doc2doc) approach (Maruf et al., 2022; Wang et al., 2023). Translation on a sentence-by-sentence basis, even with augmented contexts, remains inadequate for fully capturing coherence and consistency within texts (Fernandes et al., 2023). Various methods demonstrate the importance of context in translation; (Wu et al., 2023) process context within different components of the same encoder, while (Voita et al., 2018) use attention distributions to integrate context awareness. We are exploring a translation method that effectively bridges the gap between sentence-level and document-level approaches, aiming to harness more contextual information than sentence-level translation while avoiding the extensive resource demands of full document-level translation.
2.2 Preference Alignmnet for LLMs
Large language models (LLMs) often misalign with human preferences. Traditional approaches such as supervised fine-tuning (SFT) (Ziegler et al., 2019), reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), direct policy optimization (DPO) (Rafailov et al., 2023), and simple preference optimization (SimPO) (Meng et al., 2024) refine models post-training but demand significant computational resources. Test-time alignment offers a lightweight alternative by modifying outputs during inference without altering model parameters. Prior methods include InferAligner (Wang et al., 2024b) for cross-model guidance, ARGS (Khanov et al., 2024) with dynamic reward-based adjustment, TreeBoN (Qiu et al., 2024) leveraging tree-search in Best-of-N Sampling, and RAIN (Li et al., 2024) and URIAL (Lin et al., 2024) employing advanced prompt engineering. However, these approaches struggle when early responses are suboptimal, hindering subsequent outputs. To address this, Plan2Align introduces an MPC framework that enables predictive planning, effectively accumulating and leveraging high-quality experience to optimize response generation. By integrating structured foresight into test-time alignment, Plan2Align mitigates early-stage errors, leading to more reliable and adaptable model outputs.
Figure 2: The framework of Plan2Align. To leverage past high-quality translations for preference alignment, Plan2Align transforms the translation task into a self-rewriting process. The LLM refines its translations by revisiting previous outputs, segmenting the paragraph-level result, and aligning at the context level. A reward model, trained on preference data, selects optimal contexts to update the context buffer. Through iterative refinement, the better context buffer improves context quality and generates better rewriting inputs, achieving both paragraph-level preference alignment and context-level coherence.
3 Methodology
In this section, we introduce the Plan2Align framework (Figure 2). We describe how predictive planning enables preference alignment, the operational sequence of the framework, and how final translations are produced.
3.1 Test-Time Preference Alignment via Predictive Planning
Our key idea is to recast MT as trajectory optimization and thereby employ predictive planning for iterative translation refinement.
Machine Translation as Trajectory Optimization.
The goal of trajectory optimization is to find an optimal sequence of actions $\boldsymbol{a}^{*}=(a_{1}^{*},\cdots,a_{T}^{*})$ such that the expected trajectory-wise utility is maximized (Chua et al., 2018; Lowrey et al., 2019). More specifically: Let $s_{t}$ denote the state of the environment at the $t$-th step and let $s_{1}$ be the initial state. Consider an environment with Markov transition dynamics, where the state $s_{t+1}$ is independent of the past history given the current state $s_{t}$ and action $a_{t}$. Let $\mathcal{J}(\boldsymbol{a},\boldsymbol{s})$ denote the utility function of the sequences $\boldsymbol{a}$ and $\boldsymbol{s}$. Then, the search for $\boldsymbol{a}^{*}$ can be formulated as
$\boldsymbol{a}^{*}=\arg\max_{\boldsymbol{a}_{1:T}}\ \mathbb{E}\big{[}\mathcal{J}(\boldsymbol{a}_{1:T},\boldsymbol{s}_{1:T})\rvert s_{1}\big{]},$ (1)
where $T$ is the planning horizon and the expectation is over all the randomness in the environment.
One can interpret translation as an instance of trajectory optimization by (i) treating the source paragraph as the initial state $s_{1}$, (ii) treating each $a_{t}$ as the $t$-th segment of the translation, and (iii) setting $s_{t+1}$ as the concatenation of the source paragraph and the translations up to the $t$-th segment.
Model Predictive Control.
In general, direct optimization of (1) requires searching over all possible action sequences of length $T$ and is computationally intractable. As a predictive planning method, MPC approximately solves (1) by iteratively solving local optimization problems (Hansen et al., 2022), instead of globally optimizing the total utility in one pass. Specifically, MPC sequentially determines the action of each time step $t$ by estimating the optimal subsequence $\boldsymbol{a}^{*}_{t:t+H}$ given the current state $s_{t}$, i.e.,
$\Pi^{\text{MPC}}(s_{t})=\arg\max_{\boldsymbol{a}_{t:t+H}}\ {\mathbb{E}}\big{[}\mathcal{J}(\boldsymbol{a}_{t:t+H},\boldsymbol{s}_{t:t+H})\rvert s_{t}\big{]},$ (2)
and then executing the first action of $\Pi^{\text{MPC}}(s_{t})$. In practice, MPC solves (2) by employing a learned predictive dynamics model to generate multiple $H$-step predictive rollouts $\{(\boldsymbol{a}_{t:t+H}^{(i)},\boldsymbol{s}_{t:t+H}^{(i)})\}_{i=1}^{K}$ and obtain an approximate maximizer by taking the maximum of $\mathcal{J}(\boldsymbol{a}_{t:t+H}^{(i)},\boldsymbol{s}_{t:t+H}^{(i)})$ over the rollouts.
Predictive Planning for Preference Alignment.
In paragraph-level MT, we adopt the principle of predictive planning in MPC for preference alignment. Figure 2 illustrates the framework of Plan2Align. Specifically: (i) We train a reward model on preference data. This reward model, which plays the role of utility function in MPC, assesses translation context quality and serves as a guiding function to align paragraph-level translations with preferences. By integrating MPC, we iteratively refine translations, ensuring alignment with both linguistic quality and contextual consistency. (ii) The LLM itself can serve as a predictive dynamics model as the next state can be obtained by appending the newly generated translation segment to the current state. (iii) To enable MPC for paragraph-level MT while ensuring strong context-level performance, we further introduce two key mechanisms for accumulating and utilizing high-quality translation experiences as described in the next two subsections.
3.1.1 Context Alignment and Context Buffer
A fundamental challenge in using MPC in paragraph-level MT is: How to continuously accumulate and retain valuable translation experience? A high-scoring translated paragraph does not necessarily indicate that every part of it is of high quality. Moreover, non-extensive LLMs struggle to identify and leverage high-quality contexts from translated paragraphs due to their limited long-text comprehension capabilities.
Instead of direct MPC application in MT, we modify MPC to selectively retain high-quality contexts from multiple paragraph-level translations instead of relying on a single best trajectory. This ensures effective accumulation and utilization of translation experience. To achieve this, we establish a context buffer, which stores high-quality context extracted from multiple paragraph-level translation experiences.
[…]
Each translation is aligned through context alignment, establishing correspondences between source and translation contexts. We evaluate the aligned source and translation context pairs using a reward model trained on given preference data. Our goal is to ensure that the context buffer retains only high-quality contexts. The reward model assesses candidate context-level translations and selectively retains only the best ones. When a selected translation context surpasses a predefined buffer threshold and a matching source context already exists in the context buffer, a comparison is performed. If the new context falls below the threshold or does not surpass the quality of the existing context in the buffer, it is not stored.
Through this approach, in each iteration, the system filters high-quality contexts, continuously refining the context buffer and ensuring that subsequent translations improve over time.
3.1.2 Translations from Self-Rewriting Tasks
To leverage high-quality contexts in the context buffer for translation improvement, we frame the translation task as a self-rewriting task. To increase the diversity of translation outputs generated by the LLM, thereby enriching the context buffer with more varied contexts, we sample reference contexts from the buffer to construct the self-rewriting task input, rather than using the entire buffer content. By providing LLMs with pre-aligned content as input for self-rewriting, they can refine paragraph-level translations using past high-quality contexts while maintaining context-level coherence.
3.2 Final Translation from Context Buffer
The final translation generation differs from context buffer updates. In this phase, the stored high-quality contexts, covering nearly the entire paragraph, serve as reference translations. Unlike the self-rewriting task, which aligns individual sentences, this final translation generation stage constructs the reference translation by incorporating all relevant content from the context buffer and feeding it along with the source paragraph into the LLM. The LLM then improves the reference translation based on the source, preventing overlapping translation artifacts and ensuring fluency. Since this step involves refining a single high-quality reference translation with the source text, even non-extensive LLMs can handle the task effectively.
3.3 The Details of Context Alignment
Since the final translation relies on the quality of the context buffer, improving its quality is crucial. Therefore, continuously updating the context buffer to retain the most valuable translation experience is essential. To achieve this, context alignment must ensure proper pairing between source and translation contexts, allowing the reward model to effectively filter out low-quality translations and retain only the most reliable ones.
However, there is no well-established method to effectively extract contextual information in paragraph-level MT. To address this, we propose a simple yet effective approach for context alignment, which consists of three key steps: (i) Sentence Segmentation: We utilize the SpaCy (Honnibal et al., 2020) toolkit to segment paragraphs into a list of sentences. (ii) Sentence Alignment: We use VecAlign (Thompson and Koehn, 2020) to dynamically search for the optimal alignment between sentences. Since LLMs may introduce hallucinations, leading to sentence omissions or over-translations (resulting in extra sentences), we insert any unaligned translation sentences into their original positions and add empty strings on the source side as placeholders. (iii) Contextual Sliding Window Construction: To capture the contextual structure at the paragraph level, we adopt a sliding window approach with a window size of 3, forming the context-level aligned source-translation pairs.
zh $\rightarrow$ en zh $\rightarrow$ ru zh $\rightarrow$ de
[…]
GPT-4o 2024-08-06 ${}_{\text{~{}2024-08-06}}$ 94.58 73.06 27.77 0.10 93.74 54.20 19.89 0.00 94.54 54.55 21.93 0.00
Qwen-2.5 (14B) - 94.43 72.44 25.14 0.18 90.47 50.13 18.09 3.08 92.98 50.02 19.30 1.24
Llama-3.1 (8B) $\times$ 84.36 58.27 20.74 10.47 86.28 34.32 8.30 4.19 88.97 39.49 11.05 4.43
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 89.79 58.40 7.30 4.55 82.83 35.18 7.19 10.26 83.67 38.46 10.82 10.63
[…]
4 Experiments
4.1 Translation Dataset
Compared to sentence-level MT, paragraph-level MT has received significantly less attention, and many translation models struggle to process long-context inputs. For example, Post and Junczys-Dowmunt (2024) investigated a long-context training paradigm for machine translation but restrict the maximum sequence length to 256. ALMA Xu et al. (2024a, b) restricted their post-training datasets to a maximum sequence length of 512. To address this limitation and ensure a suitable benchmark for paragraph-level MT, we use the latest WMT’24 Discourse-Level Literary Translation benchmark (Wang et al., 2024a) for our experiments. This dataset allows for evaluating how well translation methods preserve contextual relationships during translation. The available language pairs include: Chinese →English, Chinese →German, and Chinese →Russian.
To ensure that the input length remains within the context window of our selected LLMs, we segment each instance from the training and validation sets into chunks of up to 1024 tokens. This preprocessing ensures that paragraph-level MT tasks are effectively handled within the model’s constraints. But as the readers shall see, our methodology is easy to extend to even longer context given LLM translation capability is up to par.
4.2 Preference Data and Model Training
Preference data is crucial for training alignment methods and for training the reward model used in Plan2Align. The preference data is derived from the training set of the dataset. Specifically, each instance is segmented into paragraphs of up to 1024 tokens. From each translation direction, we sample 2K paragraphs, resulting in a total of 6K paragraphs for constructing the preference data. We generate translation outputs using LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o. The translations are then evaluated using MetricX-24-XL (Juraska et al., 2024), a reference-free quality estimation (QE) model. The QE scores help us select high-quality translations: the translation with the highest score is designated as the preferred translation, while the one with the lowest score is considered not preferred. The translation with a middle score is disregarded. The resulting preference data is used to train the reward model in Plan2Align, enabling it to assess context quality. The trained reward model achieves a final accuracy of 88.53% on the validation set. Further details on the formation of preference data can be found in Appendix A.
4.3 Evaluation Metrics
We focus on optimizing LLM-based paragraph-level MT models to mitigate contextual inconsistencies, hallucinations, and omissions. However, while model-based metrics have taken over string-based metrics as the state-of-the-art, all model-based metrics (e.g. COMET) are still only trained with sentence-level data, and machine translation evaluation on paragraph-level remains an open question. A prior work Deutsch et al. (2023) argued that existing metrics generalize well to paragraph-level, but the test set in that work is a mere concatenation of existing datasets translated by sentence-level models, and does not take into account the context-level improvements we are aiming at. In fact, in our preliminary study, we observed serious pathological behaviors of these metrics when evaluating our paragraph-level outputs (see Appendix E). This calls for a significant rework of the existing evaluation paradigm.
The changes we’ve made to our machine translation evaluation paradigm are as follows:
Contextual Sliding Window.
Instead of simply feeding the source, translation and reference sentences into COMET [^1] Unbabel/wmt22-comet-da and COMET-KIWI (the reference-free version) [^2] Unbabel/wmt22-cometkiwi-da, we follow Vernikos et al. (2022) and add aligned and concatenated context as inputs to the model-based metrics. We call these metrics CW-COMET and CW-KIWI. We use the same context size of 3 as described in Section 3.3.
Sentence Segmentation and Alignment.
Similar to IWSLT evaluation paradigm for speech translation Ahmad et al. (2024), we propose to treat sentence segmentation and alignment as part of the evaluation for our paragraph-level evaluation. We follow the same general procedure as described in Section 3.3. We handle sentence alignment differently according to whether the metric requires reference input. Misaligned sentences remain in their original positions, with empty strings as placeholders where necessary. If a sliding window contains only empty strings on the one side, a default score of zero is given. This penalizes over-translation and omission, aligning with our research objectives. We verified the validity of our paradigm via a sanity check. Details are in Appendix D.
In addition to reporting model-based metrics, because over- and under-translations usually create significant misalignments across sentences, we track them by reporting the ratio of 1-0 and 0-1 alignments (1-0/0-1 Ratio), which represents the proportion of windows in the source or translation that fail to align. This is evaluated on our sentence-segmented and sentence-aligned test set. We also report document-level BLEU (d-BLEU) as some previous work has done, even it does not correlate well with human judgments Deutsch et al. (2023) and should only be treated as an auxiliary metric.
4.4 Baselines
We evaluate all training-time alignment methods on LLaMA-3.1-8B-Instruct and also use it as the backbone model for test-time alignment methods, including Plan2Align. This ensures consistency in comparing the effectiveness and aligns with our focus on small-scale LLMs. Additionally, we compare these results against the original LLaMA-3.1-8B-Instruct to further assess improvements and changes, serving as the baseline for performance. Implementation details of Plan2Align, including parameters and prompt design, are in Appendix C.
GPT-4o and Qwen-2.5-14B.
We select GPT-4o as the upper bound for performance. GPT-4o is not optimized for translation, but closely approaches that of specialized translation systems (Shahriar et al., 2024). Its precedent model, GPT-4, was the top-ranked model for 5 langauge pairs in WMT24 general machine translation shared task Kocmi et al. (2024). In addition, we also include Qwen-2.5-14B for comparison, as it demonstrates strong performance in Chinese language contexts.
Test-Time Alignment Models.
We select ARGS (Khanov et al., 2024) and RAIN (Li et al., 2024) as they both represent test-time alignment methods similar to Plan2Align but with distinct operational mechanisms. ARGS is selected for its reward-guiding method similar to Plan2Align’s; however, while Plan2Align evaluates the entire context with a reward model, ARGS dynamically recalculates scores for each token during generation. On the other hand, RAIN is selected for its ability to score newly generated tokens and perform auto-regressive self-evaluation, requiring a start-from-scratch search and backward progression. To better observe the benefits brought by MPC, we also compare it with Best-of-N and Vanilla MPC. The implementation details of test-time alignments, are available in Appendix B.1.
Training-Time Alignment Models.
We also compare Plan2Align with various training-time alignment methods. We consider SFT directly on the same preference dataset. Moreover, we include SimPO for comparison as it serves as a strong benchmark for training-time alignment in general. We also consider DPO as it serves as a mainstream representative of the training-time-alignment series, with training details available in Appendix B.2.
4.5 Quantitative Results
We present the primary results for zh→xx in Table 1. Our primary goal is not to outperform all models but rather to maximize the utility of existing models, approaching or even surpassing training-time alignment methods. Since training-time methods have direct exposure to real paragraph data, direct comparisons to test-time methods would be unfair. Instead, we emphasize whether Plan2Align can consistently improve existing model performance. We summarize the observations as follows:
Plan2Align Outperforms Existing Test-Time Alignment Methods.
Table 1 demonstrates that Plan2Align outperforms test-time alignment baselines across all translation directions. This is primarily because paragraph-level MT requires generating long outputs, and existing methods such as ARGS and RAIN guide the decoding process at the token level through a reward model. These methods struggle with error accumulation, where poor early-context translations negatively affect subsequent generations. Additionally, Table 2 highlights the efficiency of Plan2Align compared to ARGS and RAIN, detailing the time required for each approach across iterations. Since Plan2Align operates directly at the context level, it generates complete translated paragraphs without additional decoding overhead, making it both faster and more effective.
Plan2Align Closes the Performance Gap Between Test-Time and Training-Time Alignment Methods.
[…]
Best-of-N, Vanilla MPC, and Plan2Align.
As discussed in Section 1, non-extensive LLMs struggle to generate high-quality paragraph translations in a single pass. This limitation renders Best-of-N ineffective for paragraph-level MT, while directly applying MPC (i.e., Vanilla MPC) fails due to the model’s difficulty in selecting optimal contexts for subsequent translations. Table 1 shows that Best-of-N improves translation quality but does not enhance the model’s fundamental capabilities. Its effectiveness remains limited by randomness, making it an inefficient approach. Even with $N=60$ (six times the number of iterations in Plan2Align), it still underperforms compared to Plan2Align across all translation directions. Vanilla MPC achieves improvements for zh→en, but its effectiveness diminishes in zh→ru and zh→de, sometimes even falling below LLaMA-3.1’s baseline performance. This suggests that for less familiar languages, LLMs struggle to extract meaningful improvements. Plan2Align overcomes this limitation through context alignment and self-rewriting, enabling the effective application of MPC to paragraph-level MT.
AVG. Time Consumed
[…]
ARGS 1686 2343 2035
RAIN 4639 5518 4454
Plan2Align 311 313 424
Table 2: Average execution time (in seconds) per paragraph for three test-time alignment methods on the validation set. Plan2Align, with an iteration number of 3, is significantly faster than ARGS and RAIN since it avoids token-level alignment. Tests are conducted on a single NVIDIA RTX 6000 Ada Generation GPU.
[…]
Plan2Align 1 93.28 66.89 20.99 0.32
2 93.82 66.72 20.75 0.52
3 94.13 67.06 21.17 0.20
4 93.95 66.19 20.81 0.38
5 93.97 66.59 21.09 0.35
Vanilla MPC 1 89.79 58.40 17.30 4.55
2 88.00 56.26 16.45 6.42
3 83.92 53.08 15.11 10.70
4 84.66 53.60 15.26 9.93
5 84.93 54.03 15.48 9.68
Table 3: Performance of Plan2Align and Vanilla MPC across different iterations on the zh $\rightarrow$ en task.
Impact of Iteration Number in MPC.
As iterations increase, the context buffer expands, and Plan2Align selects more reference contexts. However, over-translation and omission often prevent perfect semantic alignment. Using excessive reference contexts in the self-rewriting task can overload the LLM, making it harder to extract meaningful insights due to overlapping semantics. To isolate translation difficulty effects, we focus on zh $\rightarrow$ en performance. Table 3 shows that while increasing reference contexts improves translation initially, excessive references (iteration = 4) degrade performance due to semantic ambiguity. Interestingly, Vanilla MPC exhibits an even sharper decline that is starting at iteration = 2, even in the simpler zh $\rightarrow$ en task. This supports our hypothesis that LLMs cannot directly leverage MPC effectively. Small-scale models struggle to extract useful contexts, leading to error accumulation and deteriorating translations. Plan2Align overcomes this limitation through structured context alignment and self-rewriting, ensuring stable improvements in paragraph-level MT.
5 Conclusion
In this paper, we introduced Plan2Align, a new form of model predictive planning framework for test-time preference alignment in paragraph-level machine translation. By recasting translation as a trajectory optimization problem and leveraging a self-rewriting mechanism, our approach effectively addresses key challenges such as semantic drift, omissions, and incoherence that arise in one-pass, next-token generation. Our results suggest that predictive planning can robustly accumulate and utilize high-quality translation contexts, paving the way for further improvements in long-form machine translation. Future work will explore broader language pairs and the integration of training-time strategies to further enhance translation quality. Our code will be open source under MIT license.
Limitation
One of our limitation lies in the scope of translation languages. Our experiments have focused on language pairs (e.g., Chinese to English, German, and Russian) from reproducible benchmarks, and while these choices provide valuable insights, they could potentially not encompass the full linguistic diversity present in the other translation tasks. Languages with complex morphologies, low-resource languages, or those with radically different syntactic structures may present challenges that our current formulation of context alignment and iterative refinement does not fully address. Future work has been scheduled to explore a broader set of languages to ensure that the method generalizes well and does not inadvertently favor certain language families or linguistic features over others.
Ethical and Societal Considerations
While Plan2Align is designed to enhance translation coherence and quality, its deployment raises important ethical and societal considerations. First, there is an inherent risk of bias amplification. The reward model, which drives context selection during test-time alignment, is trained on preference data that may itself contain implicit biases. If not carefully audited and diversified, this could lead to translations that systematically favor dominant linguistic norms or cultural perspectives, marginalizing underrepresented human dialects. In sum, by focusing on a limited set of languages and model architectures, there is a risk of reinforcing existing digital divides. If high-quality translation tools are predominantly developed for well-resourced languages and large-scale models, speakers of low-resource or underrepresented languages may be left behind. As researchers and practitioners, it is a long-term developing responsibility to prioritize inclusivity and fairness, ensuring that technological advances in machine translation benefit a diverse research community.
Acknowledgment
This research has been supported by NVIDIA grant and technical involvement from NVIDIA Taiwan AI R&D Center (TRDC).
Appendix A Details of Preference Data
We used MetricX-24-XL to labeled our pairwise preference datasets, the Chinese sources sentences are from WMT’24 Discourse-Level Literary Translation benchmark, we employed LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o to generate translations across three language pairs; each language pair has 2000 records, with a maximum length of 1024 tokens. The distribution of preferences, indicating the number of translation is the best translation among the three,indicating the number of translation is the best translation among the three, given that MetricX-24 scores range from 0 to 25, where 0 is the best and 25 is the worst. We removed translations scoring above 20, and if two out of three translations in a paragraph exceeded this threshold, we did not use that paragraph. The details are in Table 4. Our test-time setting, pre-trained models, and preference data will be open source under MIT license.
LLaMA Wins Gemma Wins GPT4 Wins
[…]
Table 4: The statistics of winning translations for each language pairs evaluated by MetricX-24.
Appendix B Implementation Details of Baselines
B.1 Test-Time Alignment Methods
B.1.1 ARGS
We follow the setting of ARGS, We adopt the ARGS-greedy method in ARGS as our baseline. Following the setting of ARGS-greedy. We set w to 1.5 and k to 10. For fairness, we replaced the backbone model with the same LLaMA3.1-8B-Instruct as Plan2Align, and the reward model was also replaced with the reward model used by Plan2Align.
Although ARGS settings indicate that using ARGS-greedy results in answers more closely aligned with the characteristics specified in the reward model, ARGS uses the weighted sum of logit of the token and the reward for token generation. Given that the number of tokens generated by ARGS-greedy does not exceed those produced by Plan2Align and RAIN, we included ARGS-stochastic for comparison and conducted best-of-n to optimize results, the choice of n was determined based on the average number of tokens required by Plan2Align and RAIN to generate a single translation.
However, ARGS-stochastic’s best-of-n did not surpass ARGS-greedy in performance, leading us to ultimately select ARGS-greedy as the baseline.
B.1.2 RAIN
In RAIN, we also replaced the backbone model with LLaMA3.1-8B-Instruct and replaced the self-evaluation prompt in RAIN with text in the Figure 3. For parameters in RAIN, we set value threshold V to 0.8 as the default setting of RAIN. We try four combinations of maximum and minimum number of search iterations for finding the parameter that generates the required number of tokens close to Plan2Align, which are [10,20]. The detailed configuration of the tokens generated in each parameter pairs is in Table 5.
(MinT,MaxT) AVG. tokens
[6,12] 6575
[8,16] 8972
[10,20] 13401
[12,24] 17291
Table 5: The average number of tokens required to generate a translation for RAIN in each maximum and minimum number of search iterations pairs.
Figure 3: Prompt templates for RAIN.
B.1.3 Vanilla MPC
Vanilla MPC applies MPC directly to paragraph-level machine translation (MT) by selecting the best output from the previous iteration and refining it. We conduct six iterations, with performance results presented in Table 6. The best performance for each language pair (iteration = 1, 2, 4 for zh→en, zh→ru, and zh→de, respectively) is reported in Table 1. The results show that even in the zh→en translation direction, performance continues to decline over iterations. For more challenging language pairs (zh→ru and zh→de), LLMs struggle to extract useful contexts, preventing sustained improvement and leading to unstable performance.
[…]
zh-en 1 89.79 58.40 17.30 4.55
2 88.00 56.26 16.45 6.42
3 83.92 53.08 15.11 10.70
4 84.66 53.60 15.26 9.93
5 84.93 54.03 15.48 9.68
6 84.48 52.91 15.12 10.08
zh-en 1 73.80 30.83 5.54 20.00
2 82.83 35.18 7.19 10.26
3 77.08 34.08 6.15 16.40
4 79.56 33.25 6.21 13.74
5 78.35 33.31 6.34 15.35
6 78.35 33.37 5.43 15.24
zh-en 1 83.06 39.00 10.89 11.19
2 80.88 36.74 8.95 13.03
3 79.79 36.72 9.70 14.51
4 83.67 38.46 10.82 10.63
5 82.72 38.54 10.25 11.40
6 79.35 35.87 9.54 14.87
Table 6: Performance of Vanilla MPC across different iterations.
B.2 Training-Time Alignment Methods
We choose LLaMA3.1-8B-Instruct as our experiments’ backbone model, including the reward model and training-time methods. We utilize one NVIDIA RTX 6000 Ada Generation GPU to train models with the LLaMA Factory library [^3] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024). For the training setups, the SFT model is trained on preferred data from the preference dataset, while the Reward Model, DPO, and SimPO are trained on the full preference data that includes inputs, good translations, and bad translations. All models are configured with an identical set of hyperparameters: they utilize the AdamW(Loshchilov and Hutter, 2019) optimizer and feature gradient accumulation steps fixed at 8, with a cutoff length of 2048. The maximum allowable gradient norm is restricted to 1.0 to ensure training stability. All training procedures are conducted using bf16 precision, each model undergoes a single epoch of training with a batch size of 2, and the LoRA(Hu et al., 2022) rank and LoRA alpha parameters are set to 16 across all models.
The learning rates for the models are: Rewarg Model at 1e-5, SFT at 2e-5, and both DPO and SimPO at 1e-6 as suggested by (Rafailov et al., 2023) and (Meng et al., 2024). Additionally, SimPO utilizes a beta and gamma of 2.5, which aligns with its setup, whereas SFT and DPO use a beta of 0.1 also followed by (Rafailov et al., 2023). When Plan2Align serves as a generic method, we also plan to connect the method for multi-agent (i.e., cascaded) based translation system Hu et al. (2024) in the future.
Appendix C Implementation Details of Plan2Align
In this section, we describe our parameter settings and prompt design, as well as the impact of different prompts on the final translation generation. Our full approach is outlined in Algorithm 1.
Algorithm 1 Plan2Align
0:  Source $x$, number of MPC iterations $N$, base model $\mathrm{LM}$, reward model $r$
0:  Translation result $y_{final}$
1:  Initial buffer $\text{Initial buffer }B$
2:  for $k=0\to N-2$ do
[…]
5:     for all $(w_{x},w_{y})\in(W_{x},W_{y})$ do
6:        $\mathrm{reward}\leftarrow r(w_{x},w_{y})$ // Compute rewards for each window pair
7:        Update the best window at each position to put into Buffer $B$ // Window selection
8:     end for
[…]
10:  end for
11:  return  $y_{final}$
C.1 Parameter Setting
[…]
C.2 Prompt Design
Here is the prompt configuration for generating translations in Plan2Align: Figure 4 displays our prompt template for the self-rewriting task within Plan2Align. We employ three distinct system prompts that yield three types of translations: sentence-by-sentence translation, precise translation, and imaginative translation. These prompts cater to different translation needs within the context buffer, allowing us to extract the most effective parts. Additionally, we use a context prompt to ensure the quality of the translations produced. Figures 5 and 6 showcase the prompts used for generating the final translations in Plan2Align. The distinctions between these two prompts, Reference and Annotation, are further explored in Section C.2.1.
Figure 4: Prompt template used for self-rewriting task in Plan2Align and an actual example.
Figure 5: Prompt used for generating final translation in Plan2Align (the Reference version of the prompt in Section C.2.1) and an actual example.
Figure 6: Applying the concept of the self-rewriting task to generate the final translation prompt (the Annotation version of the prompt in Section C.2.1).
C.2.1 Impact of Final Translation Task
In Plan2Align, final translation generation differs from context buffer updates. The final translation is produced by assembling buffer-stored contexts in source sentence order, forming a reference translation (Reference). Conversely, during context buffer updates, translated sentences are annotated beneath their corresponding source sentences (Annotation). Table 7 shows that the Reference approach outperforms Annotation in zh→ru and zh→de. However, for zh→en, the Annotation method generally performs better. This difference arises because annotating each source sentence with context buffer information and refining segments independently, although not requiring explicit alignment, demands higher semantic understanding. This makes it effective for zh→en but burdensome for the other two translation directions. In contrast, the Reference approach offers greater flexibility in refining the entire paragraph.
[…]
zh-en Annotation 94.02 67.83 22.16 0.38
Reference 94.13 67.06 21.17 0.20
zh-ru Annotation 80.58 35.42 10.09 12.83
Reference 91.23 40.11 10.69 1.49
zh-de Annotation 80.25 37.27 9.09 13.54
Reference 92.60 44.62 10.69 1.44
Table 7: Performance of Plan2Align across two different final translation tasks.
Appendix D Evaluation Sanity Check
To verify the validity of our sentence segmentation and alignment approach in the context of machine translation evaluation, we conduct an evaluation using the test set for WMT24 general machine translation translation shared task, as it is a sentence-aligned dataset. We select Zh-En and En-Zh translation data for our sanity check. Specifically, we concatenate sentences with spaces to simulate paragraph structures while ensuring that each paragraph does not exceed 1024 tokens, resulting in a total of 1181 paragraphs. In addition, we simulate imperfect alignments with under- and over-translations by creating two extra scenarios: (1) randomly dropping 10% of source sentences and (2) randomly dropping 10% of target sentences.
For simplicity, we do not generate system translation and directly evaluate the source-reference pair using KIWI-22. For our evaluation paradigm to be valid, the score distribution should change similarly when over- and under-translations are introduced. We also track the change in sentence pairs in our sanity check, but we expect that to change with dropped sentences because one-to-many and many-to-one alignments may be generated.
Table 8 presents the alignment performance under different conditions. It can be seen that our score distribution reacts similarly to the score calculated with the ground-truth alignment when under- and over-translation errors are introduced. This demonstrates that our evaluation method effectively captures semantic alignment and validates the applicability of CW-COMET and CW-KIWI in reflecting context-level translation quality.
Scenario # Pairs Mean Std.
Perfectly Aligned 22956 78.03 7.77
Our Alignment 22310 77.07 9.79
Drop 10% (Source) 22956 73.68 15.20
Our Alignment 21222 73.81 14.20
Drop 10% (Target) 22956 75.25 11.78
Our Alignment 21433 74.97 11.94
Table 8: Alignment performance across different conditions. The change in our score maintains high consistency with the score calculated with perfect alignment.
Appendix E MetricX Scoring Failure Cases
Here are three failure cases for metricX-24, each illustrated with different models and language pairs:
RAIN Model, zh-ru Pair.
The translation output appears nonsensical (see Figure 7), yet metricX-24 awarded it a score of 8.76, indicating a seemingly better translation under a metric where lower scores denote superior results.
ARGS Model, zh-de Pair.
As shown in Figure 8, ARGS produced repetitive phrases without completing the translation from the original zh-de text, reaching the max token limit. Despite this, it received a score of 8.47 from metricX-24.
LLaMA3.1 Model, zh-en Pair.
As shown in Figure 9, the translation did not start from the beginning of the text, and the meaning conveyed was incomplete. Nevertheless, metricX-24 gave it a high score of 9.85, which inaccurately suggests a less effective translation.
Figure 7: Case 1: Failed evaluation on RAIN, zh-ru, metricX score: 8.76, produced nonsensical output.
Figure 8: Case 2: Failed evaluation on ARGS, zh-de, metricX-qe: 8.47, produced incomplete translation and repeated words and sentences
Figure 9: Case 3: Failed evaluation on LLaMA3.1-8B, zh-en, metricX-qe: 9.85, did not start the translation from the beginning and the meaning is incomplete.