### v2

# Plan2Align: Predictive Planning Based Test-Time Preference Alignment for Large Language Models

Abstract
Aligning Large Language Models with Preference Fine-Tuning is often resource-intensive. Test-time alignment techniques that do not modify the underlying model—such as prompting and guided decoding—offer a lightweight alternative. However, existing test-time alignment methods primarily improve short responses and fail to ensure coherence over extended contexts due to the myopic nature of token-level alignment. Moreover, these methods often incur a slowdown during inference. To address these challenges, we propose Plan2Align, a test-time alignment framework that formulates text generation as a predictive planning problem. Plan2Align adapts Model Predictive Control (MPC) to iteratively refine output by rolling out multiple complete responses and optimizing each segment. To more rigorously evaluate the effectiveness and efficiency, we focus on the more challenging task of long-text generation. Experiments on the long-form response subset of the HH-RLHF dataset and the WMT’24 Discourse-Level Literary Translation demonstrate that Plan2Align significantly enhances the performance of base LLMs. Compared to existing training-time and test-time alignment methods on LLaMA-3.1 8B, Plan2Align achieves comparable or superior results, while also delivering improved inference efficiency relative to prior test-time alignment approaches.
footnotetext: Equal contribution. Correspondence to Kuang-Da Wang <gdwang.cs10@nycu.edu.tw>, Ping-Chun Hsieh <pinghsieh@nycu.edu.tw>, and Chao-Han Huck Yang <hucky@nvidia.com>.
1 Introduction
Figure 1: Plan2Align mirrors control rollouts, selecting optimal segments for alignment.
The emergence of Large Language Models (LLMs), such as the GPT series (Brown et al., 2020; Achiam et al., 2023), LLaMAs (Touvron et al., 2023a, b), and Gemma (Team et al., 2024), has demonstrated remarkable efficacy in a wide range of NLP tasks (Stiennon et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2023; Yu et al., 2024; Zhong et al., 2024). While these models exhibit strong performance out of the box, aligning their outputs to human preferences remains critical, especially for smaller-scale LLMs. For instance, in machine translation (Alves et al., 2024), smaller LLMs (e.g., under 10B parameters) frequently suffer from hallucinations, omissions, and semantic drift (Wu et al., 2024). Thus, aligning LLM outputs to preferences remains an essential yet challenging problem. Methods such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) and Direct Preference Optimization (DPO) (Rafailov et al., 2023) have shown promising results in aligning preferences. However, these methods are typically resource-intensive.
Consequently, there is growing interest in test-time alignment techniques, which operate without modifying model parameters through strategies like prompting (Lin et al., 2024) or guided decoding (Khanov et al., 2024; Wang et al., 2024b; Li et al., 2024; Kong et al., 2024). However, existing methods face two fundamental challenges:
  1. 1.
    Limited Adaptivity: Many methods rely on guided decoding or value estimation, predicting outputs without the ability to revise earlier errors. This makes them brittle when the generation drifts from preferred trajectories, as they lack mechanisms for correction.
  2. 2.
    High Inference Overhead: Due to their focus on token or short-span alignment, these methods frequently introduce significant inference overhead, limiting their efficiency.
To address these challenges, we propose Plan2Align, a novel test-time replanning framework that formulates preference alignment as a predictive planning problem. Inspired by the principle of Model Predictive Control (MPC), a classic predictive planning method for trajectory optimization (Rao, 2014) widely used in robotics (Tassa et al., 2012, 2014; Nagabandi et al., 2020) and motion planning (Ji et al., 2016). As a predictive planning method, MPC optimizes a sequence of actions by continuously updating future plans based on past observations (Camacho and Bordons, 2007; Kouvaritakis and Cannon, 2016). Analogously, Plan2Align treats text generation as a trajectory in which each segment of output informs the next through an iterative feedback process. This framing enables the model to revise responses at the segment level [^1] The definition of a segment may vary depending on the task. For instance, in long-text translation, a segment may consist of three sentences, whereas in HH-RLHF, segments are defined by response position: the first third, middle third, and final third. The core idea is to leverage informative segments to guide the generation of more coherent and aligned outputs., improving both consistency and efficiency over long-form outputs. Figure 1 illustrates this correspondence, highlighting the intuition behind our approach.
Specifically, we introduce two core mechanisms:
  1. 1.
    Model-Predictive Segment Selection: We extend the MPC principle to selectively retain high-quality segments from multiple full responses using a reference buffer, rather than relying on a single best trajectory. This enables more effective accumulation and utilization of alignment-relevant signals over the course of generation.
  2. 2.
    Rewrite-Guided Replanning: To better guide small-scale LLMs toward aligned responses, we provide richer signals by recasting the task as a rewriting task. This structured prompting mechanism enables each iteration to build upon prior high-quality segments—not limited to the initial prompt but drawn from different locations in the response. By grounding future generation in selected high-quality text spans, our approach reduces reasoning difficulty for the model and promotes output consistency.
In this manner, Plan2Align enables efficient alignment and content refinement over a limited number of iterations, offering a practical and effective solution to preference alignment in small-scale LLMs.
Our contributions are summarized as follows:
  • We introduce a novel test-time alignment framework, transforming the alignment task into a predictive planning problem leveraging MPC principles, thus ensuring effective alignment.
  • We propose Plan2Align, integrating Model-Predictive Segment Selection and Iterative Replanning Generation, enabling efficient incremental improvements of generated text through effective utilization of high-quality segments within limited iterations.
  • We empirically evaluate Plan2Align using LLaMA-3.1 8B on challenging datasets including the WMT’24 discourse-level literary translation Benchmark and the long-form response subset of HH-RLHF. Our method demonstrates substantial improvements, achieving results comparable or superior to training-time alignment methods and outperforming existing test-time alignment approaches.
2 Related Work
2.1 Preference Alignment through Fine-tuning
Aligning large language models (LLMs) with human preferences has traditionally relied on post-training strategies. Supervised fine-tuning (SFT) (Ziegler et al., 2019) and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) are widely used but computationally expensive. Direct Preference Optimization (DPO) (Rafailov et al., 2023) simplifies RLHF by converting it into a supervised learning objective, though it requires managing dual policies. More recent approaches like SimPO (Meng et al., 2024) and Contrastive Preference Optimization (CPO) (Xu et al., 2024) reduce memory and resource demands using reference models and contrastive signals. Despite these improvements, fine-tuning methods remain rigid and slow to adapt to changing data or objectives, posing challenges in dynamic environments.
2.2 Test-Time Preference Alignment
Test-time alignment offers a lightweight and adaptable alternative to fine-tuning by guiding generation during inference without updating model parameters. Beyond prompting and in-context learning, guided decoding is a major category of test-time alignment methods that explicitly control the generation process using external signals. ARGS (Khanov et al., 2024) is a representative example that incorporates reward model guidance at the token level, and InferAligner (Wang et al., 2024b) adopts a similar strategy. Among guided decoding methods, there are also approaches that directly modify internal representations. For instance, RE-Control (Kong et al., 2024) trains a value function on hidden states using the Bellman equation, and applies gradient-based optimization to align with preferences. TreeBoN (Qiu et al., 2024) and RAIN (Li et al., 2024) leverage tree-based structures: TreeBoN combines tree search with Best-of-N sampling, while RAIN performs self-evaluation without relying on a reward model to align preferences.
While these methods enhance alignment, they often incur high latency and computational overhead due to token-level control, especially in long-form generation. Moreover, value-guided decoding approaches struggle with inaccurate value estimation and instability when generation drifts from preferred trajectories. We propose a test-time alignment method that balances efficiency and alignment quality, achieving robust performance even in long-form generation settings.
3 Methodology
3.1 Test-Time Preference Alignment via Predictive Planning
Our key idea is to recast preference alignment as trajectory optimization and thereby employ predictive planning for iterative refinement.
Text Generation Optimization as Trajectory Optimization.
The goal of trajectory optimization is to find an optimal sequence of actions $\boldsymbol{a}^{*}=(a_{1}^{*},\cdots,a_{T}^{*})$ such that the expected trajectory-wise utility is maximized (Chua et al., 2018; Lowrey et al., 2019). More specifically: Let $s_{t}$ denote the state of the environment at the $t$-th step and let $s_{1}$ be the initial state. Consider an environment with Markov transition dynamics, where the state $s_{t+1}$ is independent of the past history given the current state $s_{t}$ and action $a_{t}$. Let $\mathcal{J}(\boldsymbol{a},\boldsymbol{s})$ denote the utility function of the sequences $\boldsymbol{a}$ and $\boldsymbol{s}$. Then, the search for $\boldsymbol{a}^{*}$ can be formulated as
$\boldsymbol{a}^{*}=\arg\max_{\boldsymbol{a}_{1:T}}\ \mathbb{E}\big{[}\mathcal{J}(\boldsymbol{a}_{1:T},\boldsymbol{s}_{1:T})\rvert s_{1}\big{]},$ (1)
where $T$ is the planning horizon and the expectation is over all the randomness in the environment.
One can interpret text generation as an instance of trajectory optimization by (i) treating the input prompt as the initial state $s_{1}$, (ii) treating each $a_{t}$ as the $t$-th segment of the response, and (iii) setting $s_{t+1}$ as the concatenation of the input prompt and the response up to the $t$-th segment.
Model Predictive Control.
In general, direct optimization of (1) requires searching over all possible action sequences of length $T$ and is computationally intractable. As a predictive planning method, MPC approximately solves (1) by iteratively solving local optimization problems (Hansen et al., 2022), instead of globally optimizing the total utility in one pass. Specifically, MPC sequentially determines the action of each time step $t$ by estimating the optimal subsequence $\boldsymbol{a}^{*}_{t:t+H}$ given the current state $s_{t}$, i.e.,
$\Pi^{\text{MPC}}(s_{t})=\arg\max_{\boldsymbol{a}_{t:t+H}}\ {\mathbb{E}}\big{[}\mathcal{J}(\boldsymbol{a}_{t:t+H},\boldsymbol{s}_{t:t+H})\rvert s_{t}\big{]},$ (2)
and then executing the first action of $\Pi^{\text{MPC}}(s_{t})$. In practice, MPC solves (2) by employing a learned predictive dynamics model to generate multiple $H$-step predictive rollouts $\{(\boldsymbol{a}_{t:t+H}^{(i)},\boldsymbol{s}_{t:t+H}^{(i)})\}_{i=1}^{K}$ and obtain an approximate maximizer by taking the maximum of $\mathcal{J}(\boldsymbol{a}_{t:t+H}^{(i)},\boldsymbol{s}_{t:t+H}^{(i)})$ over the rollouts.
Model Predictive Path Integral (MPPI).
MPPI (Williams et al., 2015) is a stochastic, sampling-based instantiation of model predictive control that reformulates the optimization in Eq. (2) as a path integral over trajectories. Rather than relying on gradient-based updates, MPPI approximates the control distribution by generating multiple candidate trajectories with random perturbations and computing their expected utility. Given a predictive model of environment dynamics, MPPI samples $K$ control sequences $\{(\boldsymbol{a}_{t:t+H}^{(i)}\}_{i=1}^{K}$ from a proposal distribution (typically Gaussian) and simulates the corresponding state trajectories. Each trajectory is scored by a cost or reward function $\mathcal{J}$, and the final control is computed as a weighted average:
$\boldsymbol{a}_{t}=\frac{\sum_{i=1}^{K}\exp\left(-\frac{1}{\lambda}\mathcal{J}^{(i)}\right)\boldsymbol{a}_{t}^{(i)}}{\sum_{i=1}^{K}\exp\left(-\frac{1}{\lambda}\mathcal{J}^{(i)}\right)},$ (3)
where $\lambda$ is a temperature parameter controlling the trade-off between exploration and exploitation. Unlike deterministic MPC, MPPI enables smoother updates by aggregating over multiple high-reward rollouts rather than relying on a single best trajectory.
Figure 2: Overview of Plan2Align. Left: Plan2Align adapts MPPI control to language generation by evaluating response segments with a reward model and reusing high-quality ones from a reference buffer. Right: We implement this with Model-Predictive Segment Selection (purple), where segment-level rollouts are scored and stored in a buffer, and Rewrite-Guided Replanning (gray), which uses aligned high-quality segments as rewriting task to progressively improve output.
Adapting MPPI for Preference Alignment.
Plan2Align draws inspiration from MPPI control and reinterprets its core principles for test-time generation with LLMs. As shown in Figure 2, we align components of MPPI with analogous operations in LLM-based generation:
  • In MPPI, a policy and world model are used to sample multiple future trajectories, each representing a sequence of actions. In Plan2Align, the LLM plays both roles, generating diverse candidate responses conditioned on an input prompt. These responses serve as discrete trajectories in language space.
  • Instead of evaluating entire trajectories, Plan2Align decomposes responses into segments and performs fine-grained segment-level evaluation using a reward model trained on preference data. This segment-wise assessment replaces MPPI’s scalar cost for each rollout, allowing more localized credit assignment and reuse of high-quality spans.
  • Rather than selecting one best trajectory or optimizing a control policy, Plan2Align stores high-scoring segments into a reference buffer. This buffer is used to construct rewriting prompts, enabling the LLM to adjust its behavior over successive iterations without modifying model parameters.
This adaptation retains the predictive planning spirit of MPPI—roll out, evaluate, and replan—but translates it into a discrete, segment-level alignment framework suited for long-form generation. As shown in the right half of Figure 2, rewriting prompts are constructed from a reference buffer of high-quality segments, allowing the model to iteratively improve its outputs during inference. The mechanisms that support this loop are described in the next two subsections.
3.2 Model-Predictive Segment Selection for Managing High-Quality Experience
We adapt the MPPI paradigm by selectively retaining high-quality segments from multiple responses, regardless of their position, instead of relying solely on the leading segment of the best-scoring output. This approach enables more efficient accumulation of useful generation trajectories. We implement this using a reference buffer that stores top-scoring segments across responses. Given that segment positions are fixed within a response, we maintain a buffer of size $k$ for each position, preserving the top-$k$ segments to guide subsequent replanning steps.
3.3 Rewrite-Guided Replanning for Improving Outputs
While high-quality experience can be retained through the reference buffer, effectively leveraging it to guide LLMs toward preference-aligned outputs remains a non-trivial challenge. This is particularly true for small-scale LLMs, which have limited capacity for contextual reasoning. Reducing the complexity of the replanning process is therefore critical.
To address this, we reframe the replanning task as a rewriting task. By providing LLMs with pre-aligned segments as input, the model can refine responses using high-quality content from previous iterations. Before rewriting, the system filters and retains only those segments that exceed a predefined quality threshold [^2] The threshold is typically based on the reward score of a preferred response. If the threshold is set too high, buffer updates may occur too slowly. If set too low, the buffer may include suboptimal segments.. This ensures that only informative and preference-aligned segments are carried forward, preventing contamination from low-quality content.
In this way, the model is not required to infer the exact insertion points for each segment. Instead, segments are positioned at fixed locations, and their combination serves as the basis for rewriting. This design lowers the reasoning burden while still allowing the model to produce improved responses. These refined responses are then used to update the buffer in the next iteration, promoting stable improvements in generation quality and alignment with preferences.
3.3.1 Accelerating the Preference Alignment Process
To accelerate preference alignment, Plan2Align aims to minimize the number of rewriting iterations by efficiently discovering diverse high-quality segments. We introduce stochasticity into rewriting task construction to promote exploration and increase the likelihood of retrieving useful content. Below, we describe how segments are defined and sampled for two different tasks. The corresponding rewriting prompt templates for each task are provided in Appendix G.1.2 and Appendix G.2.2.
Response Generation.
Response generation offers a relatively simple setting. The LLM generates $n$ responses per iteration, and each is divided into $m$ sentence-level segments of equal length. Segment positions are fixed to simplify alignment. Because the reward model assigns a single score per response, we uniformly propagate the score to all its segments. To construct rewriting prompts, we sample one segment per position from the top-$k$ highest scoring segments in the buffer. This randomness encourages diversity and helps discover better segments. If a sampled segment falls below a predefined quality threshold, we omit it, prompting the model to refine that portion.
Translation.
In translation, segment alignment is naturally guided by source–target correspondence. We define source-based segments as three-sentence spans with overlap via a sliding window (e.g., sentences 1–3, 2–4). Unlike response generation, the reward model can independently evaluate each segment, enabling more precise quality assessment and a reference buffer that better captures high-quality translations. Translation introduces additional challenges due to alignment noise between source and target segments, which can degrade prompt quality. To address this, we sample a subset of segments during intermediate iterations and progressively increase the sample size. The output translation uses all buffered segments. See Appendix Algorithm 1 for details.
4 Experiments
We evaluate Plan2Align on two challenging long-form generation tasks—long-form response generation and paragraph-level machine translation—to better reveal performance differences.
4.1 Preference Dataset
4.1.1 Long-Form Response Dataset
For long-form response generation, we use the Dahoas/full-hh-rlhf [^3] https://huggingface.co/datasets/Dahoas/full-hh-rlhf dataset, which is widely adopted for LLM alignment. This dataset is designed to improve AI assistant behavior in terms of helpfulness and harmlessness. Each sample consists of a prompt and two responses, with one labeled as preferred based on human judgments. Since the response lengths in the dataset vary significantly, we select samples based on the length of the chosen responses. Specifically, we construct the training set using the top 6K samples with the longest chosen responses from training set, and using the top 1024 longest chosen responses from the testing set to construct test set.
4.1.2 Paragraph-Level Translation Dataset
To establish a suitable preference dataset for long-text MT, we use the latest WMT’24 Discourse-Level Literary Translation benchmark (Wang et al., 2024a) for our experiments. This dataset allows for evaluating how well translation methods preserve contextual relationships during translation. The available language pairs include: Chinese →English, Chinese →German, and Chinese →Russian. To keep inputs within the context window of our chosen LLMs, we segment each training and validation instance into chunks of up to 1024 tokens, using GPT-4’s tokenizer to standardize token length. This preprocessing ensures that paragraph-level MT tasks are handled within the model’s constraints.
Preference Data Generation.
The preference data is derived from the training set of the dataset. Specifically, each instance is segmented into paragraphs of up to 1024 tokens. From each translation direction, we sample 2K paragraphs, resulting in a total of 6K paragraphs for constructing the preference data. We generate translation outputs using LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o. The translations are then evaluated using MetricX-24-XL (Juraska et al., 2024), a reference-free quality estimation (QE) model. The QE scores help us select high-quality translations: the translation with the highest score is designated as the preferred translation, while the one with the lowest score is considered not preferred. The translation with a middle score is disregarded.
4.2 Reward Model Training
Preference data is crucial for training alignment methods and for training the reward model used in Plan2Align. The collected preference data is utilized to train the reward model in Plan2Align, allowing it to evaluate the quality of contextual responses. The reward model trained for long-form responses achieves a validation accuracy of 83.78%, while the model trained for paragraph-level translation attains a validation accuracy of 88.53%. Further details on the formation of preference data can be found in Appendix C, and detail of training can be found in Appendix F.2.
4.3 Evaluation Metrics
4.3.1 Long-Form Responses
Drawing on prior work (Khanov et al., 2024; Kong et al., 2024), we evaluate responses quality using the following five metrics:
Average Reward reflects the average score assigned by the reward model during Plan2Align and ARGS inference, which indicates alignment with helpfulness and harmlessness preferences. Coherence is measured as the cosine similarity between the embeddings of the prompt and its continuation, computed using a SimCSE model (Gao et al., 2021), capturing how well the generated response maintains topical consistency. Diversity assesses lexical variation in the output, calculated by aggregating n-gram repetition rates. For a given continuation $y$, the diversity score is defined as $\text{diversity}(y)=\prod_{n=2}^{4}\frac{\text{unique $n$-grams}(y)}{\text{total $n$-grams}(y)}$, with higher values indicating broader vocabulary use. Win Rate captures the proportion of pairwise comparisons in which a model’s response is rated higher than a reference response by GPT-4 (OpenAI et al., 2024). Following the ARGS evaluation protocol (Khanov et al., 2024), GPT-4 is prompted to evaluate overall response quality, considering helpfulness, harmlessness, relevance, accuracy, depth, creativity, and detail. The full evaluation prompt is included in Appendix E. Lastly, Inference Time refers to the average time (in seconds) required to generate a single response, reflecting the computational efficiency of each method.
4.3.2 Paragraph-Level MT
In machine translation tasks, most evaluations rely on existing model-based metrics. Following CPO (Xu et al., 2024), we adopt COMET [^4] Unbabel/wmt22-comet-da and COMET-KIWI [^5] Unbabel/wmt22-cometkiwi-da (the reference-free version) for our evaluation. While model-based metrics have taken over string-based metrics as the state-of-the-art, all model-based metrics (e.g. COMET) are still only trained with sentence-level data, and machine translation evaluation on paragraph-level remains an open question. A prior work Deutsch et al. (2023) argued that existing metrics generalize well to paragraph-level, but the test set in that work is a mere concatenation of existing datasets translated by sentence-level models, and does not take into account the context-level improvements we are aiming at. In fact, in our preliminary study, we observed serious pathological behaviors of these metrics when evaluating our paragraph-level outputs (see Appendix H.2). This calls for a significant rework of the existing evaluation paradigm. The changes we’ve made to our machine translation evaluation paradigm are as follows:
Contextual Sliding Window.
Instead of simply feeding the source, translation and reference sentences into COMET and COMET-KIWI, we follow Vernikos et al. (2022) and add aligned and concatenated context as inputs to the model-based metrics. We call these metrics CW-COMET and CW-KIWI. We use the same size of 3 as described in Section 3.3.1.
Sentence Segmentation and Alignment.
Similar to IWSLT evaluation paradigm for speech translation Ahmad et al. (2024), we propose to treat sentence segmentation and alignment as part of the evaluation for our paragraph-level evaluation. We follow the same general procedure as described in Section 3.3.1. We handle sentence alignment differently according to whether the metric requires reference input. Misaligned sentences remain in their original positions, with empty strings as placeholders where necessary. If a sliding window contains only empty strings on the one side, a default score of zero is given. This penalizes over-translation and omission, aligning with our research objectives. We verified the validity of our paradigm via a sanity check. Details are in Appendix H.
4.4 Baselines
To ensures consistency in comparing the effectiveness, we evaluate all training-time alignment methods on LLaMA-3.1-8B-Instruct and also use it as the backbone model for test-time alignment methods, including Plan2Align. Additionally, we compare these results against the original LLaMA-3.1-8B-Instruct to further assess improvements and changes, serving as the baseline for performance. Implementation details of Plan2Align, including parameters and prompt design, are in Appendix G.
Test-Time Alignment Models.
We select ARGS (Khanov et al., 2024) and RAIN (Li et al., 2024) as they both represent test-time alignment methods similar to Plan2Align but with distinct operational mechanisms. To better isolate the effect of predictive planning, we also compare Plan2Align with Best-of-N sampling. Additionally, we introduce a simplified variant of our method, referred to as Vanilla MPC. This ablation removes both the segment-level evaluation and the use of a reference buffer. Instead of constructing rewriting tasks, Vanilla MPC simply selects the best-scoring response or translation from each iteration and feeds it into the next generation round for refinement. Implementation details of all test-time alignment baselines are provided in Appendix F.1.
Training-Time Alignment Models.
We also compare Plan2Align with training-time alignment methods. We consider SFT directly on the same preference dataset, which is typically a strong baseline for translation tasks. Moreover, we include SimPO for comparison as it serves as a strong benchmark for training-time alignment in general. We also consider DPO as it serves as a mainstream representative of the training-time-alignment series, with training details available in Appendix F.2.
GPT-4o and Qwen-2.5-14B.
For paragraph-level MT, we select GPT-4o as the upper bound for performance. GPT-4o is not optimized for translation, but closely approaches that of specialized translation systems (Shahriar et al., 2024). We also include Qwen-2.5-14B, as it has demonstrated strong performance in Chinese language tasks and serves as a representative open-source alternative.
[…]
Llama-3.1 (8B) $\times$ 2.9528 0.6160 0.7936
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 4.0776 0.6364 0.7885
[…]
Figure 3: Results on the test set of Dahoas/full-hh-rlhf. Left: Quantitative comparison across base models, training-time, and test-time alignment methods. For the two groups showing results from alignment methods, best numbers in respective groups are boldfaced, and best numbers among all alignment methods are also underlined. Our method are highlighted. Right: Win rate of Plan2Align responses compared to DPO and Best-of-20.
4.5 Quantitative Results
4.5.1 Results on Long-Form Responses
We present the primary results for long-form response generation in Figure 3. This task provides a challenging benchmark for evaluating how well different alignment methods can enhance the generation quality of base model. We summarize the observations as follows:
Superior to training-time and test-time baselines.
Compared to strong baselines such as DPO and SimPO, our method achieves higher average reward, coherence, and diversity, all without any model fine-tuning. Moreover, it delivers a substantial improvement over the base model. We further assess response quality using GPT-4 preference judgments on the 1,024 test samples. The reward model trained with Plan2Align achieves an accuracy of 75.71% on this evaluation set, demonstrating strong alignment with GPT-4 preferences. To provide a comparative analysis, we evaluate Plan2Align against the best-performing methods from both training-time and test-time alignment: DPO and Best-of-20, respectively. As illustrated in the right side of Figure 3, Plan2Align significantly outperforms both baselines in head-to-head comparisons. GPT-4’s win/loss/tie statistics for each comparison are detailed in the figure
Effective alignment with fewer rollouts.
Our method performs 3 iterations with 3 rollouts each, in addition to the initial LLM output, totaling 10 generations. In contrast, Best-of-20 samples twice as many outputs, yet still underperforms relative to Plan2Align. This demonstrates that our gains are not attributable to random sampling. Even under equal generation cost (Best-of-10), Plan2Align maintains a clear advantage, validating the benefit of planning and rewriting over naive sampling. Despite its improved alignment quality, Plan2Align remains efficient at inference time. While the baseline LLaMA3.1-8B model requires 8 seconds to generate a single response, ARGS takes 363 seconds (44×), and RAIN takes 1930 seconds (242×). In contrast, Plan2Align completes inference in 108 seconds—only 13× the cost of vanilla generation—while being substantially more efficient than other test-time alignment methods. These results highlight that Plan2Align achieves a strong balance between alignment performance and generation efficiency.
Ablation via Vanilla MPC.
In this variant, we remove segment-level evaluation and the reference buffer, selecting only the best full response at each iteration for refinement. While this strategy still improves upon the base model, it lags behind both Plan2Align and Best-of-10, highlighting the necessity of our core design choices for effective preference alignment.
zh $\rightarrow$ en zh $\rightarrow$ ru zh $\rightarrow$ de
[…]
GPT-4o 2024-08-06 ${}_{\text{~{}2024-08-06}}$ 94.58 73.06 93.74 54.20 94.54 54.55
Qwen-2.5 (14B) - 94.43 72.44 90.47 50.13 92.98 50.02
Llama-3.1 (8B) $\times$ 84.36 58.27 86.28 34.32 88.97 39.49
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 88.50 57.06 73.84 31.45 90.32 42.20
[…]
4.5.2 Results on Paragraph-Level MT
We present the primary results for zh→xx in Table 1. Compared to HH-RLHF, we consider long-text translation a more challenging task, especially for small-scale LLMs. This is reflected in Table 1, where we observe a large gap between the base model and SoTA models. Therefore, our primary goal is not to outperform all models but rather to maximize the utility of existing models, approaching or even surpassing training-time alignment methods. We summarize the observations as follows:
Plan2Align Narrows the Gap Between Test-Time and Training-Time Alignment.
As shown in Table 1, Plan2Align consistently outperforms all existing test-time alignment baselines across translation directions. Unlike token-level methods such as ARGS and RAIN, which guide decoding with fine-grained reward signals and suffer from early error accumulation, Plan2Align operates at the segment level and incorporates high-quality partial translation into a rewriting framework. This enables greater coherence and contextual consistency in paragraph-level translation. Moreover, Plan2Align narrows the gap with training-time alignment methods, despite lacking the direct exposure to paragraph-level data that fine-tuning approaches typically benefit from. For zh→en, Plan2Align outperforms all training-time baselines under the CW-COMET metric, which is generally considered more reliable due to its reference-based nature. It even surpasses the GPT-4o and Qwen-2.5, highlighting the effectiveness of test-time planning in aligning long-form translations with human preferences. For zh→ru, although CW-COMET shows Plan2Align achieving strong performance, CW-KIWI scores remain lower than those of training-time methods. This discrepancy is likely due to limitations in the KIWI-22 evaluation model for this language pair, which fails to fully capture Plan2Align’s improvements. For zh→de, Plan2Align does not outperform training-time approaches, but it delivers comparable results. This suggests that while our method is generally effective, its performance may still be limited by the base model’s inherent translation abilities.
Best-of-N, Vanilla MPC, and Plan2Align.
Table 1 shows that Best-of-N sampling remains limited in effectiveness due to its reliance on unguided randomness. Even with $N=60$ (5 times the number of rollouts in Plan2Align), it still underperforms compared to Plan2Align across all translation directions. Similarly, the Vanilla MPC variant also fails to match the performance of Plan2Align. Without segment-level evaluation or rewriting, it lacks the structural guidance necessary for effective refinement. These results further validate the design of Plan2Align and its ability to perform efficient and robust preference alignment across domains.
5 Conclusion, Limitations and Future Work
We introduced Plan2Align, a test-time model predictive planning framework for preference alignment. By treating generation as trajectory optimization, Plan2Align incrementally refines outputs through segment-level evaluation and rewriting, leading to consistent improvements without modifying model parameters. While effective, Plan2Align depends on the base model’s expressiveness—it can only amplify behaviors already latent in the model. When desirable outputs lie far from the model’s distribution, test-time planning alone may be insufficient. Future work will explore how iterative distribution shifts arise under Plan2Align and how these mechanisms can be integrated with training-time alignment to support more robust generalization.
Appendix A Limitations and Future Work
While Plan2Align effectively improves output quality through test-time preference alignment, it is fundamentally constrained by the expressiveness of the underlying language model. It can only steer generation toward outputs already supported by the model’s capabilities, and may struggle when the desired outputs lie far from the model’s original distribution. In future work, we aim to better understand the conditions under which iterative alignment shifts the output distribution effectively. This could enable stronger integration with training-time methods, such as lightweight fine-tuning or reward model co-optimization, to improve robustness in harder settings and under distribution shifts.
Appendix B Broader Impacts
We anticipate that Plan2Align will contribute positively to the development of more controllable and aligned language models, particularly in low-resource or safety-critical applications where fine-tuning is impractical. By enabling preference-aware generation at test time without modifying model weights, our method may lower the barrier to deploying LLMs in real-world scenarios. However, test-time alignment methods also introduce new risks, such as reinforcing spurious patterns if reward signals are poorly calibrated. To mitigate this, we recommend careful design and auditing of reward models, especially in domains involving sensitive or subjective outputs.
Appendix C Details of Preference Data
Both of two reward model serves as the alignment objective for the Plan2Align framework.
C.1 Long-Form Responses
For the long-form response task, we construct a pairwise preference dataset using the Dahoas/full-hh-rlhf dataset, where each pair consists of a chosen and a rejected response. These pairs are used as preference data to train the reward model. To ensure a fair evaluation and maintain consistency in response length, we filter both the training and test sets to include only samples with continuation lengths close to 1024 tokens. Specifically, we select 6,000 training samples that fall within this length range.
C.2 Paragraph-Level MT
For paragraph-Level machine translation task, we used MetricX-24-XL to labeled our pairwise preference datasets, the Chinese sources sentences are from WMT’24 Discourse-Level Literary Translation benchmark, we employed LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o to generate translations across three language pairs; each language pair has 2000 records, with a maximum length of 1024 tokens. The distribution of preferences, indicating the number of translation is the best translation among the three,indicating the number of translation is the best translation among the three, given that MetricX-24 scores range from 0 to 25, where 0 is the best and 25 is the worst. We removed translations scoring above 20, and if two out of three translations in a paragraph exceeded this threshold, we did not use that paragraph. The details are in Table 2. Our test-time setting, pre-trained models, and preference data will be open source under MIT license.
LLaMA Wins Gemma Wins GPT4 Wins
[…]
Table 2: The statistics of winning translations for each language pairs evaluated by MetricX-24.
Appendix D Supplementary Results
Since the results in the main paper are condensed versions with showing only a subset of the metrics evaluated, we attach the full breakdown of the results in Table 3 for readers’ reference. In addition to reporting model-based metrics, because over- and under-translations usually create significant misalignments across sentences, we track them by reporting the ratio of 1-0 and 0-1 alignments (1-0/0-1 Ratio), which represents the proportion of windows in the source or translation that fail to align. We also report document-level BLEU (d-BLEU) as some previous work has done, even it does not correlate well with human judgments (7) and should only be treated as an auxiliary metric.
zh $\rightarrow$ en zh $\rightarrow$ ru zh $\rightarrow$ de
[…]
GPT-4o 2024-08-06 ${}_{\text{~{}2024-08-06}}$ 94.58 73.06 27.77 0.10 93.74 54.20 19.89 0.00 94.54 54.55 21.93 0.00
Qwen-2.5 (14B) - 94.43 72.44 25.14 0.18 90.47 50.13 18.09 3.08 92.98 50.02 19.30 1.24
Llama-3.1 (8B) $\times$ 84.36 58.27 20.74 10.47 86.28 34.32 8.30 4.19 88.97 39.49 11.05 4.43
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 88.50 57.06 13.39 6.04 73.84 31.45 3.96 19.55 90.32 42.20 10.19 2.91
[…]
Appendix E Evaluation prompt of GPT-4
Figure 4 is the prompt template for the GPT-4 evaluation, we follow the ARGS evaluation setup (15), with a slight modification:Our prompt explicitly instructs GPT-4 to consider how well each assistant addresses the user’s question, rather than to evaluate the response in isolation. The prompt asks GPT-4 to rate responses based on overall performance, considering criteria such as helpfulness, harmlessness, relevance, accuracy, depth, creativity, and level of detail
Figure 4: System prompt and user prompt of GPT-4 evaluation.
Appendix F Implementation Details of Baselines
F.1 Test-Time Alignment Methods
F.1.1 ARGS
We follow the setting of ARGS, We adopt the ARGS-greedy method in ARGS as our baseline. Following the setting of ARGS-greedy. We set w to 1.5 and k to 10. For fairness, we replaced the backbone model with the same LLaMA3.1-8B-Instruct as Plan2Align, and the reward model was also replaced with the reward model used by Plan2Align.
Although ARGS settings indicate that using ARGS-greedy results in answers more closely aligned with the characteristics specified in the reward model, ARGS uses the weighted sum of logit of the token and the reward for token generation. Given that the number of tokens generated by ARGS-greedy does not exceed those produced by Plan2Align and RAIN, we included ARGS-stochastic for comparison and conducted best-of-n to optimize results, the choice of n was determined based on the average number of tokens required by Plan2Align and RAIN to generate a single translation.
However, ARGS-stochastic’s best-of-n did not surpass ARGS-greedy in performance, leading us to ultimately select ARGS-greedy as the baseline.
F.1.2 RAIN
In RAIN, we also replaced the backbone model with LLaMA3.1-8B-Instruct and replaced the self-evaluation prompt in RAIN with text in the Figure 5. For parameters in RAIN, we set value threshold V to 0.8 as the default setting of RAIN. We try four combinations of maximum and minimum number of search iterations for finding the parameter that generates the required number of tokens close to Plan2Align, which are [10,20]. The detailed configuration of the tokens generated in each parameter pairs is in Table 4.
(MinT,MaxT) AVG. tokens
[6,12] 6575
[8,16] 8972
[10,20] 13401
[12,24] 17291
Table 4: The average number of tokens required to generate a translation for RAIN in each maximum and minimum number of search iterations pairs.
Figure 5: Prompt templates for RAIN.
F.2 Training-Time Alignment Methods and Reward Model
We adopt LLaMA3.1-8B-Instruct as the backbone model for all experiments, including the reward model and training-time alignment methods. Model training is conducted using a single NVIDIA RTX 6000 Ada Generation GPU, with the implementation based on the LLaMA Factory library [^6] https://github.com/hiyouga/LLaMA-Factory (44).
For training setups, the SFT model is trained solely on the preferred responses from the preference dataset, while the Reward Model, DPO, and SimPO are trained on the full preference data, which includes the input, chosen, and rejected responses, as detailed in Appendix C.
All models are trained using identical hyperparameters: the AdamW optimizer (20) with gradient accumulation steps set to 8, a sequence cutoff length of 2048 tokens, and a maximum gradient norm clipped at 1.0 for stability. Training is performed using bf16 precision, with a batch size of 2, for one epoch. We apply LoRA (12) with a rank of 16 and alpha of 128 across all models.
The learning rates are configured as follows: 2e-5 for the Reward Model and SFT, and 5e-6 for both DPO and SimPO, in accordance with recommendations from Rafailov et al. (27) and Meng et al. (22). Additionally, SimPO is trained with a gamma value of 0.4, while DPO uses a beta value of 0.1, consistent with the original settings proposed in their respective works.
Appendix G Implementation Details of Plan2Align and Vanilla MPC
In this section, we describe our parameter settings and prompt design. Our full approach is outlined in Algorithm 1.
Algorithm 1 Plan2Align: Test-Time Preference Alignment via Predictive Planning
0:  Input prompt $x$, base LLM $\mathcal{M}$, reward model $\mathcal{R}$, number of iterations $T$, rollouts per iteration $K$
0:  Final response $y^{\text{final}}$
1:  Initialize reference buffer $\mathcal{B}\leftarrow\emptyset$
2:  Initialize candidate response set $\mathcal{Y}\leftarrow\emptyset$
3:  Generate initial response $y^{(0)}\leftarrow\mathcal{M}(x)$
4:  Add $(y^{(0)},\mathcal{R}(y^{(0)}))$ to $\mathcal{Y}$
5:  Add segments from $y^{(0)}$ to buffer $\mathcal{B}$
6:  for $t=1$ to $T$ do
7:     Sample $K$ candidate responses $\{y_{k}^{(t)}\}_{k=1}^{K}\leftarrow\mathcal{M}(x,\mathcal{B})$Rollouts}
8:     for each $y_{k}^{(t)}$ do
9:        Compute reward $r_{k}\leftarrow\mathcal{R}(y_{k}^{(t)})$
10:        if task is not translation then
11:           Add $(y_{k}^{(t)},r_{k})$ to $\mathcal{Y}$
12:        end if
13:        Decompose $y_{k}^{(t)}$ into segments $\{s_{j}\}$
14:        for each segment $s_{j}$ do
15:           Compute segment reward $r_{j}\leftarrow\mathcal{R}(s_{j})$
16:           if $r_{j}>$ threshold then
17:              Add $s_{j}$ to buffer $\mathcal{B}$ at position $j$
18:           end if
19:        end for
20:     end for
21:     if task is translation then
22:        Construct final rewriting prompt using all segments in $\mathcal{B}$
23:        Generate final translation $y^{\text{trans}}\leftarrow\mathcal{M}(x,\mathcal{B}_{\text{all}})$
24:        Compute reward $r^{\text{trans}}\leftarrow\mathcal{R}(y^{\text{trans}})$
25:        Add $(y^{\text{trans}},r^{\text{trans}})$ to $\mathcal{Y}$
26:     end if
27:  end for
28:  Select $y^{\text{final}}\leftarrow\arg\max_{(y,r)\in\mathcal{Y}}r$
29:  return  $y^{\text{final}}$
G.1 Long-Form Responses Generation
G.1.1 Parameter Setting
To generate diverse translations for comparison, we design three distinct system prompts for the rewriting task, resulting in three translation outputs per iteration. The segment size is fixed at 3 sentences, and the number of iterations is set to 3. The reference buffer threshold is fixed at 3.
G.1.2 Rewriting Prompt Design
We present the prompt configuration used for generating translations in Plan2Align. As shown in Figure 8, our prompt template is designed for the rewriting task within Plan2Align. We utilize three distinct system prompts, each aimed at a specific type of improvement: making the text more concise and clear, more detailed and informative, or more balanced and well-structured.
G.2 Paragraph-Level Machine Translation
G.2.1 Parameter Setting
To generate diverse translations for comparison, we design three distinct system prompts for the rewriting task, resulting in 3 translation outputs per iteration. The segment size is fixed at 3 sentences, and the number of iterations is set to 3. The reference buffer threshold is fixed at 1, as the translation task already introduces randomness through sampling at different segment positions. More specifically, when constructing the rewriting input from the reference buffer, we adopt a sampling strategy to encourage diversity and uncover more valuable segments. As the number of iterations increases, the reference buffer accumulates more content. To leverage this growing pool, we adaptively increase the number of reference segments used by Plan2Align. In particular, we define the number of sampled segments as a function of the current iteration using the formula: sampled segment number = current iteration number + 5.
G.2.2 Rewriting Prompt Design
We present the prompt configuration for generating translations in Plan2Align: Figure 6 displays our prompt template for the rewriting task within Plan2Align. We employ three distinct system prompts that yield three types of translations: sentence-by-sentence translation, precise translation, and imaginative translation. These prompts cater to different translation needs within the reference buffer, allowing us to extract the most effective parts. Figures 7 showcase the prompts used for generating the output translations in Plan2Align.
G.3 Implementation of Vanilla MPC
Vanilla MPC serves as a simplified ablation of Plan2Align, aiming to test whether repeated refinement alone—without segment-level planning—is sufficient to improve generation quality. At each iteration, the model performs multiple rollouts and selects the response or translation with the highest reward score, as determined by the reward model. This top-performing output is then passed to the language model as input for further refinement. After several iterations, the final output is selected from all responses based on reward scores.
We present the full algorithmic steps of Vanilla MPC in Appendix Algorithm 2.
System Prompts for Response Generation.
To encourage meaningful revisions in long-form response generation, we define three distinct system prompts. For fairness, they are closely aligned with those used in Plan2Align. As shown in Figure 8, these prompts are designed to guide the model toward generating improved outputs with varying qualities.
  • Improve the given response. Make it more concise and clear. Respond only with the improved answer.
  • Improve the given response. Make it more detailed and informative. Respond only with the improved answer.
  • Improve the given response. Make it balanced and well-structured. Respond only with the improved answer.
System Prompts for Paragraph-Level Translation.
To guide stylistic refinement in paragraph-level translation, we define three distinct system prompts, each corresponding to a different translation intent. As shown in Figure 6, these prompts encourage the model to revise previous translations with specific goals in mind.
  • You are a meticulous translator. Provide a literal, word-for-word translation that preserves the structure and meaning of each individual word.
  • You are a professional translator. Deliver a clear, formal, and precise translation that faithfully conveys the original meaning.
  • You are a creative and expressive translator. Render the text in a vivid and imaginative way, as if narrating a captivating story.
In addition, we provide a context prompt that presents the previously translated result:
  • In the previous evaluation, here is the result you have translated before: history_mt. Please consider this result and improve it to give a better translation.
This setup encourages the model to generate improved translations by leveraging prior outputs, guided by both explicit intent and contextual information.
Algorithm 2 Vanilla MPC for Test-Time Refinement
0:  Input $x$, base LLM $\mathcal{M}$, reward model $\mathcal{R}$, number of iterations $T$, rollouts per iteration $K$
0:  Final response $y^{\text{final}}$
1:  Initialize candidate set $\mathcal{Y}\leftarrow\emptyset$
2:  Generate initial response $y^{(0)}\leftarrow\mathcal{M}(x)$
3:  Add $(y^{(0)},\mathcal{R}(y^{(0)}))$ to $\mathcal{Y}$
4:  for $t=1$ to $T$ do
5:     Generate $K$ rollouts $\{y_{k}^{(t)}\}_{k=1}^{K}\leftarrow\mathcal{M}(y^{(t-1)},\text{prompt}_{t})$
6:     for each $y_{k}^{(t)}$ do
7:        Compute reward $r_{k}\leftarrow\mathcal{R}(y_{k}^{(t)})$
8:        Add $(y_{k}^{(t)},r_{k})$ to $\mathcal{Y}$
9:     end for
10:     Set $y^{(t)}\leftarrow\arg\max_{y}\mathcal{R}(y)$ over current $K$ rollouts
11:  end for
12:  return $y^{\text{final}}\leftarrow\arg\max_{(y,r)\in\mathcal{Y}}r$
Figure 6: The template of rewriting prompt used for translation task in Plan2Align and an actual example.
Figure 7: Prompt used for generating final translation in Plan2Align and an actual example.
Figure 8: The template of rewriting prompt used for response generation task in Plan2Align and an actual example.
Appendix H Evaluation Sanity Check
To verify the validity of our sentence segmentation and alignment approach in the context of machine translation evaluation, we conduct an evaluation using the test set for WMT24 general machine translation translation shared task, as it is a sentence-aligned dataset. We select Zh-En and En-Zh translation data for our sanity check. Specifically, we concatenate sentences with spaces to simulate paragraph structures while ensuring that each paragraph does not exceed 1024 tokens, resulting in a total of 1181 paragraphs. In addition, we simulate imperfect alignments with under- and over-translations by creating two extra scenarios: (1) randomly dropping 10% of source sentences and (2) randomly dropping 10% of target sentences.
For simplicity, we do not generate system translation and directly evaluate the source-reference pair using KIWI-22. For our evaluation paradigm to be valid, the score distribution should change similarly when over- and under-translations are introduced. We also track the change in sentence pairs in our sanity check, but we expect that to change with dropped sentences because one-to-many and many-to-one alignments may be generated.
Table 5 presents the alignment performance under different conditions. It can be seen that our score distribution reacts similarly to the score calculated with the ground-truth alignment when under- and over-translation errors are introduced. This demonstrates that our evaluation method effectively captures semantic alignment and validates the applicability of CW-COMET and CW-KIWI in reflecting context-level translation quality.
Scenario # Pairs Mean Std.
Perfectly Aligned 22956 78.03 7.77
Our Alignment 22310 77.07 9.79
Drop 10% (Source) 22956 73.68 15.20
Our Alignment 21222 73.81 14.20
Drop 10% (Target) 22956 75.25 11.78
Our Alignment 21433 74.97 11.94
Table 5: Alignment performance across different conditions. The change in our score maintains high consistency with the score calculated with perfect alignment.
H.1 Limitations of MetricX in Detecting Translation Errors
We conduct a preliminary study to assess whether state-of-the-art MT metrics can reliably detect under- and over-translation errors when applied to concatenated sentence-level outputs. Using MetricX-24 and its reference-free variant MetricX-24-QE from the WMT 2024 Metrics Shared Task (8), we evaluate three language pairs: en-de, en-es, and ja-zh. To stay within the 1024-token input limit of MetricX-24, we filter out overly long examples.
We simulate under-translation by removing sentences from the target side, and over-translation by removing sentences from both source and reference. We limit ourselves to removing at most two sentences since removing more would leave too few documents for reliable Kendall’s $\tau$ correlation calculations. We measure the performance of the metrics by computing Kendall’s $\tau$ correlations between the metric outputs and human evaluation scores. For cases with introduced under-/over- translate, we assign $25$ as the human-annotated MQM score for each , which is then converted into z-score in accordance with each human annotator’s scoring distribution. Finally we average the segment-level human judgment scores as the document-level scores.
As shown in Figure 9, both metrics exhibit limited sensitivity to these common error types. These findings motivate our use of CW-COMET and CW-KIWI, which explicitly account for cross-sentence coherence and alignment in long-form translation evaluation.
Figure 9: The Kendall’s $\tau$ correlations of MetricX-24 and MetricX-24-QE show limited sensitivity to over- and under-translation; sentences with more than three drops are insufficient to estimate correlations reliably.
H.2 MetricX Scoring Failure Cases
In addition to its limited ability to detect translation errors, we also find that MetricX-24 occasionally produces nonsensical judgments. Below, we present three failure cases from different models and language pairs to illustrate this issue.
RAIN Model, zh-ru Pair.
The translation output appears nonsensical (see Figure 10), yet metricX-24 awarded it a score of 8.76, indicating a seemingly better translation under a metric where lower scores denote superior results.
ARGS Model, zh-de Pair.
As shown in Figure 11, ARGS produced repetitive phrases without completing the translation from the original zh-de text, reaching the max token limit. Despite this, it received a score of 8.47 from metricX-24.
LLaMA3.1 Model, zh-en Pair.
As shown in Figure 12, the translation did not start from the beginning of the text, and the meaning conveyed was incomplete. Nevertheless, metricX-24 gave it a high score of 9.85, which inaccurately suggests a less effective translation.
Figure 10: Case 1: Failed evaluation on RAIN, zh-ru, metricX score: 8.76, produced nonsensical output.
Figure 11: Case 2: Failed evaluation on ARGS, zh-de, metricX-qe: 8.47, produced incomplete translation and repeated words and sentences
Figure 12: Case 3: Failed evaluation on LLaMA3.1-8B, zh-en, metricX-qe: 9.85, did not start the translation from the beginning and the meaning is incomplete.